# Efficient Multi-agent Navigation with Lightweight DRL Policy

### v1

@@ 4.1 Network structure
We first only use LSTM to encode the LiDAR observation. The policy is trained in a static environment and works well in static and 6-agent test environments. However, it performs poorly when the robot’s number increases over 6. Thus, we add an attention module to overcome the issue. As shown in Figure 3, we use a three-layer LSTM module to encode the laser measurements $o_{s}^{t}$, with both hidden sizes = 256, and get $S$. Then, $K$ and $V$ are calculated by two different FC layers by using $S$, in which the output sizes = 256. At the same time, $[o_{g}^{t},\ o_{v}^{t}]$ passes through two FC layers, which output sizes = 256, to obtain $Q$. Then, the encoding of $o_{s}^{t}$, $X$, is obtained by multi-head attention:
@@
where $d_{k}$ is the size of $K$. The final encoding is obtained by the weighted sum of $X$ and $S$ with an adaptive residual adjustment, $\alpha X+(1-\alpha)S$, where $\alpha$ updates with the policy and the value network. In actor and critic, the final encoding is decoded by two fully connected layer (FC) layers with output size = 128 and output size = 2 to get the mean of the policy $[v^{t},\omega^{t}]$, and another two FC layers with output size = 128 and output size = 1 to get the value network $V_{\phi}(o^{t})$. Note that each LSTM layer is followed by a rectified linear unit (ReLU), and so is the FC layer. Additionally, the last FC layer of the actor and critic modules has no ReLU layers.
4.2 Simulation platform
To bridge the gap between simulation and the real world, we propose a multi-agent training platform based on a physics-based simulator such as Pybullet (Coumans and Bai (2016–2021)) and Gazebo. In most 2-D simulators, the actions of the agent are usually perfectly executed, and their positions are usually calculated by $x^{t+1}=x^{t}+v^{t}\Delta t+\frac{1}{2}a^{t}\Delta t^{2}$. In a physics-based simulator, the agent state is computed by the physics engine, i.e., $x^{t+1}=T(x^{t},v^{t},a^{t},f^{t},...)$, where $T$ is the transfer function of the physics engine, $f^{t}$ is the force acting on the agent. The transfer function $T$ makes the execution of the action randomness. A straightforward example is when the straight action executes in different types of simulators, the results differ. As shown in Figure 4. The 2-D simulator executes the action without bias, while the agent in the physics-based simulator deviates from the straight line, and the deviation is greater in real-world robots. This indicates that training on a physics-based simulator may help the agent learn more realistic policies, thereby facilitating the deployment of the policy on actual robots. To better train the policy, we add Gaussian noise to the agent’s perception to increase randomness further.
(a) 2-D simulator.
(b) 3-D simulator.
(c) Real robot.
We propose a local replay mechanism for the trajectory collection of the agent’s collision aiming to speed up the collection of multi-agent interaction trajectories. When the agent collides, the agent will be reset back to the state before the $N$ simulation steps so that the agent can quickly accumulate the collision data and speed up the training, where $N=300$ in this article. However, when the local replay mechanism is deployed in a multi-agent scenario, the reset agent may appear in front of the normal agent, causing the normal agent to be disturbed. Sudden disturbances may make the policy conservative. However, we found that in some cases, agents perform slowing or stopping actions to wait for other agents to leave.
Besides, to make the simulator generalize, we adopt the architecture of Gym (Brockman et al. (2016)), and we have diligently maintained the interfaces intending to provide users with a more convenient experience when utilizing the simulator. The simulator supports the random addition of random or queued agents or obstacles and controls the movement of each agent through a action set $\mathcal{A}$. At the end of a simulator step, the simulator returns a observed state set $\mathcal{O}$. The update frequency of the simulator can be modified to meet the policy training at different control frequencies. All of these functionalities are designed for easy reconfiguration. For example, to accommodate users seeking to evaluate Safe Reinforcement Learning (Safe-RL, Achiam and Amodei (2019)), we have reconfigured our custom-designed simulator to include a constrained version by changing two functions of the origin version.
4.3 Training process
We treat distributed multi-agent navigation as plenty of single agents navigating in an unknown environment, meaning each agent collects the trajectory respectively. As shown in Figure 3, the training process can be divided into trajectory collection and policy update. During the trajectory collection, each agent first observes the LiDAR $o_{s}^{t}$, target point $o_{g}^{t}$, and velocity $o_{v}^{t}$. After that, it utilizes the policy and value network to calculate a value $v^{t}$ and an action mean $[v^{t},\omega^{t}]$. From these, an action $a^{t}$ is sampled based on the policy distribution $\pi_{\theta}(a^{t}|o^{t})$ and executed. The simulator then responds with a reward $r^{t}$, the next state $o^{t+1}$, and a termination signal $done^{t}$. The tuple of $<o^{t},a^{t},r^{t},o^{t+1},done^{t},V_{\phi}(o^{t})>$ is stored in replay buffer. This process is repeated until a specified number of iterations have been completed. Subsequently, the trajectories stored in the relay buffer are used to calculate the total loss $L^{Total}$ for updating the policy and value networks. After updating the networks, the replay buffer will be cleared, and the agents collect trajectories again. The two stages alternate until convergence or episode limit is reached.
Alg 1 shows a more detailed training process. In lines 3-9, each agent collects the trajectories respectively and calculates the values using the value network under the same environment. After specific steps, in lines 10-15, according to the trajectories collected by each agent, the corresponding advantages of each trajectory are calculated through General Advantage Estimation (GAE, Schulman et al. (2015)). Then, we merge and shuffle the trajectories to get training data. In lines 17-24, we calculate the policy loss by $L^{{PPO}}=\text{min}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)}A,\text{clip}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)},1-\epsilon,1+\epsilon)A)$, the value loss by $L^{V}=\max((V_{\theta}-V_{targ})^{2},(\operatorname{clip}(V_{\theta},V_{\theta_{old}}-\varepsilon,V_{\theta_{old}}+\varepsilon)-V_{tart})^{2})$, where $V_{targ}=A+V_{\phi}(s)$ and the entropy loss by $L^{{Entropy}}=H(\pi_{\theta}(a^{t}|o^{t}))$. We get the total loss by weighted summing $L^{{Total}}=-L^{{PPO}}+\alpha L^{V}+\beta L^{{Entopy}}$, where $\alpha$ and $\beta$ are hyperparameters. We use the Adam Optimizer (Kingma and Ba (2014)) to update the policy and value network w.r.t $L^{Total}$. The algorithm proceeds to the next iteration until a specified iteration is reached. We also use some tricks to make the training more stable. The tricks include reward scaling (Gao et al. (2022)), mini-batch updating, advantage normalizing, and learning rate decay.
We use two scenarios during training: single-agent and multi-agent. As we clip the distance input to the network, the size of the scenario does not affect the training. Thus, in the single-agent scenario, the initial point and the goal point of navigation are 10 m apart and the multi-agent scenario’s initial and goal navigation points are 16 m apart. The two scenarios are shown in Figure 5.
(a) Single-agent scenario.
(b) Multi-agent scenario.
@@ 4.3 Training process
Table 1: Experiment with 1 robot and 15 obstacles.
success rate(%) collision rate(%) trap rate (%) average step
IJRR (Fan et al. (2020)) 91.3 8.6 0.1 1383.23
  Linear policy 67.2 2.4 30.4 550.40
LSTM policy 93.1 6.9 0.0 1187.73
LSTM+Attention 95.9 3.9 0.2 1300.78
  (Safe) LSTM+Attention 97.6 2 0.04 1353.63
Table 2: Experiment with 6 robots and 5 obstacles.
success rate(%) collision rate(%) trap rate (%) average step
NH-ORCA (Alonso-Mora et al. (2010)) 98.4 0.33 1.27 1198.82
IJRR (Fan et al. (2020)) 85.03 14.63 0.33 1617.84
LSTM policy 91.89 8.15 0.0 981.01
LSTM+Attention 94.78 4.68 0.53 1805.94
Figure 6: Multi-agent test without obstacle.
The measurement dimension $L_{n}$ of the LiDAR of the policy is different, and the max length of the LiDAR beam is 4 m. We first train the policy with $L_{n}=36$, which works successfully in the single-agent scenario. In the multi-agent scenario, the $L_{n}$ of LSTM policy and LSTM+Attention are 72. Because, if $L_{n}=36$, the agent may lose information when other agents stay in the internal of the LiDAR beam, with the biggest internal = 0.34 m, while $L_{n}=72$, the biggest internal = 0.17 m is enough for most situations. The field of view (FOV) is set to $\pi$ rad. Additionally, we stack the LiDAR observations of 5 consecutive frames to provide more comprehensive observation information. The agent’s velocity is set within the $[0,1]$ m/s, and the angular velocity is set within the $[-\pi,\pi]$ rad/s. The control frequency used in training is 60 Hz, which can make the reward variation of each action more apparent and maintain efficient control. Finally, the maximum number of simulation steps is set to 2500, sufficient for the agent to finish navigation through the optimal path.
5 Experiments
To evaluate the ability of the policy, we test it in both simulation and the real world compared with NH-ORCA (Alonso-Mora et al. (2010)) and the policy of Fan et al. (2020). In the following section, we refer to the policy of Fan et al. (2020) by CNN-policy, and the measurement dimension of the LiDAR $L_{n}=512$. The LSTM+Attention policy is the proposed policy. We conducted an ablation experiment by removing the attention mechanism from the proposed policy, which is LSTM policy. Additionally, we replaced the LSTM layer with a Linear layer and get the Linear policy, in order to validate the effectiveness of the LSTM layer. In the simulation tests, we use some metrics to evaluate the policy. The metrics consist of the success rate, the collision rate, the trap rate, and the average step. The success rate means the agent’s success rate navigates to the goal in time without collision. The collision rate means the rate at which the agent collides with other objects. The trap rate means the rate at which agent cannot reach the goal in time and without collision. The average step is the average simulation step of the agent to reach the goal successfully. We use the simulation step rather than the cost time of navigation because the cost time depends on the calculation speed of the simulator, which cannot exactly reflect the cost time of navigation. Finally, all simulation tests are performed 1000 times.
5.1 Simulation test
[…]
@@ 5.1 Simulation test
Figure 8: Real-world single-agent experiment. The turtlebot3 first passes through narrow passages, then crosses through the obstacles, and finally reaches the goal point. The turtlebot3 reaches the goal point in 46 s.
5.2 Real-world test
We use turtlebot3 to test the policy’s performance in the real world. As the max linear velocity and angular velocity of the turtlebot3 are 0.2 m/s and 1 rad/s, we scale the output of our policy to fit the ability of turtlebot3. Besides, the control frequency of turtlebot3 is set to 10 Hz. The observation of the LiDAR given by turtlebot3 with parameters are FOV $=2\pi$ rad, and LiDAR beam number = 360. We only use 72 LiDAR beam with FOV $=0.8\pi$ rad. We align the rest of the policy with the simulation settings. We use the wheel odometry and LiDAR sensor on the turtlebot3 for perception. Wheel odometry calculates the robot’s relative position and angle to the goal point, while the LiDAR sensor is employed to gather LiDAR observation. Although it is possible to let each turtlebot3 calculate its action independently, it is unable to install the corresponding library, like Pytorch Paszke et al. (2017), on the onboard Raspberry Pi. Therefore, we control turtlebot3 through the Robot Operating System (ROS, Quigley et al. (2009)) communication protocol by using a host computer.
In the single-agent scenario, we test the performance of the policy through narrow passages. The turtlebot3 passes and avoids obstacles with unknown positions and shapes to reach the goal point. As shown in Figure 8, the turtlebot3 starts at a room and passes through a passage to get in the room with the goal point in 22 s. Then, the turtlebot3 crosses the obstacles to reach the goal point in 24 s. At last, the turtlebot3 finishes the navigation when approaching the goal point.
In the multi-agent scenario, we test the policy’s performance in avoiding other agents and sudden disturbances. A turtlebot3 needs to go through obstacles and avoid other turtlebot3. Besides, a human who wants to interfere with turtlebot3 is a sudden disturbance. As shown in Figure 7, at first, the human appears in front of a turtlebot3 to force it to change its trajectory, and after that, the human directly crosses through two turtlebot3. Then, the human continues to interfere with other turtlebot3 until it reaches its goal point. All turtlebot3s successfully avoid collision with other turtlebot3s and the human.
These two experiments prove that the policies trained on our proposed simulation platform can be directly deployed on real robots, with only adjustments made to the values of output actions based on the real robot’s capabilities.
6 Conclusion
This paper proposes a lightweight, end-to-end multi-agent navigation policy that translates raw LiDAR observation directly into action commands, eliminating the need to distinguish between dynamic and static obstacles and enabling navigation through narrow gaps. A physics-engine-based simulation platform is introduced for multi-agent training, bridging the gap between simulation and reality. Our method, trained using RL, demonstrates robust performance in simulated and real-world environments, even in malicious obstructions.
Key contributions include the proposal of a distributed navigation policy trained in stochastic environments, the utilization of LSTM to mitigate information loss in POMDP, the development of a multi-agent training platform to facilitate idea validation and safe RL, and the introduction of a new reward function to address suboptimal action selection. The policy’s performance is evaluated in simulation and real-world settings, highlighting its effectiveness and potential for future enhancements.
In the future, we will explore extending the policy to more complex environments, incorporating passing through traffic, or expanding the work to Unmanned aerial vehicles (UAVs). We will further refine the reward function and the policy to enhance performance in edge cases, like U-shaped obstacles.
{acks}
This work is supported by Shenzhen Science and Technology Program under Grant RCBS20221008093305007, Grant 20231115141459001, Young Elite Scientists Sponsorship Program by CAST under Grant 2023QNRC001, and High level of special funds (G03034K003) from Southern University of Science and Technology, Shenzhen, China.

### v3

@@ III-E Network Structure
As shown in Fig. 2, we use a three-layer LSTM module to encode the laser measurements $o_{s}^{t}$ with both hidden sizes = 256 and get LSTM encoding $S$. Then, $K$ and $V$ are computed using two FC layers, each with the output sizes = 256, applied to the encoding $S$. At the same time, $[o_{g}^{t},\ o_{v}^{t}]$ passes through two FC layers, which output sizes = 256, to obtain $Q$. Then, the encoding of the LiDAR observations, $X$, is obtained by multi-head attention:
@@
where $d_{k}$ is the size of $K$. The final encoding is obtained by the weighted sum of $X$ and $S$ with an adaptive residual adjustment, $\alpha X+(1-\alpha)S$, where $\alpha$ updates with the network. In actor and critic, the final encoding is decoded by two FC layers with output size = 128 and output size = 2 to get the mean of the policy $[v^{t},\omega^{t}]$, and another two FC layers with output size = 128 and output size = 1 to get the value network $V_{\phi}(o^{t})$. Note that each LSTM layer is followed by a rectified linear unit (ReLU), and so is the FC layer. Additionally, the last FC layer of the actor and critic modules has no ReLU layers.
IV EXPERIMENT EVALUATION
IV-A Experimental Setup
All models are implemented with PYTORCH on a server with a GTX4090 GPU and CPU of Intel Xeon 8336C. We test the model and select their optimal parameter settings.
To evaluate the ability of the policy, we test it in both simulation and the real world compared with NH-ORCA [6] and the policy of [12]. In the following section, we refer to the policy of [12] by the CNN policy and the measurement dimension of the LiDAR $L_{n}=512$. The LSTM-Attention policy is the proposed policy. We conduct an ablation experiment by removing the attention mechanism from the proposed policy, which is the LSTM policy. Additionally, we replace the LSTM layer with the Linear layer and obtain the Linear policy to validate the effectiveness of the LSTM layer.
In the simulation tests, we use four metrics to evaluate the policy. The metrics consist of the success rate, the collision rate, the trap rate, and the average step. The success rate means the agent’s success rate navigates to the goal in time without collision. The collision rate means the rate at which the agent collides with other objects. The trap rate means the rate at which agent cannot reach the goal in time and without collision. The average step is the average simulation step of the agent to reach the goal successfully. We use simulation steps instead of navigation cost time as the metric because the latter depends on the computational speed of the simulator, which may not accurately reflect real-world navigation time. Finally, we perform all simulation tests 1000 times.
Figure 6: Multi-agent test without obstacle.
IV-B Simulation Test
@@
Figure 7: Real-world single-agent experiment. The TurtleBot3 first passes through narrow passages, crosses through the obstacles, and finally reaches the goal point. The TurtleBot3 reaches the goal point in 46 s.
IV-C Real-World Test
To evaluate the policy’s performance in real-world scenarios, we deploy it on a TurtleBot3 platform. The TurtleBot3’s maximum linear and angular velocities are limited to 0.2 m/s and 1 rad/s, respectively; thus, we scale the policy’s output to align with these constraints. Additionally, the control frequency is set to 10 Hz. The LiDAR sensor on the TrutleBot3 provides FOV $=2\pi$ rad with 360 beams, but we only utilize 72 beams within a reduced FOV $=0.8\pi$ rad to match our experimental requirements. We align the rest of the policy with the simulation settings. We use the wheel odometry and LiDAR sensor on the TurtleBot3 for perception. Wheel odometry calculates the robot’s relative position and orientation to the goal, while the LiDAR sensor captures the surrounding environment observations.
In the single-agent scenario, we evaluate the policy’s performance by navigating the TurtleBot3 through narrow passages. The TurtleBot3 successfully passes the passages and avoids obstacles of unknown positions and shapes while reaching the goal point. As illustrated in Fig. 7, the TurtleBot3 begins in a room, traverses a narrow passage to enter the goal room in 22 s, and then avoids obstacles to achieve the goal point in an additional 24 s. Finally, the TurtleBot3 completes the navigation task as it approaches the goal point.
In the multi-agent scenario, we evaluate the policy’s performance in avoiding other agents and responding to sudden disturbances. Each TurtleBot3 needs to navigate through obstacles while avoiding the other robots and human. In this evaluation, the human acts as a dynamic disturbance by intentionally interfering with the TurtleBot3s. As shown in Fig. 8, the human first appears in front of a TurtleBot3 to force it to alter its trajectory, then crosses directly between two TurtleBot3s. Subsequently, the human continues interfering with one TurtleBot3 until it reaches its goal point. Despite these disturbances, all TurtleBot3s successfully avoid collisions with each other and the human.
The navigation algorithm demonstrates robust performance in both single-agent and multi-agent scenarios. In the single-agent scenario, the TurtleBot3 successfully navigates through narrow passages and avoids obstacles of unknown positions and shapes, efficiently reaching the goal point. In multi-agent scenario, the algorithm enables TurtleBot3s to avoid collisions with other robots and respond to dynamic disturbances, such as intentional human interference. The TurtleBot3s dynamically adjust their trajectories to accommodate sudden obstacles while maintaining collision-free navigation. These results highlight the algorithm’s adaptability and reliability in handling static and dynamic challenges, ensuring safe and efficient navigation in real-world applications.
@@ IV-C Real-World Test
V CONCLUSION AND FUTURE WORK
We present a lightweight, end-to-end multi-agent navigation framework that directly maps raw LiDAR observations to action commands, eliminating the need for explicit obstacle classification while enabling efficient navigation through constrained spaces. Our approach leverages a physics-enhanced simulation platform to bridge the simulation-to-real gap, facilitating practical multi-agent training. The reinforcement learning-based solution demonstrates robust performance across simulated and real-world environments, including challenging scenarios with adversarial obstructions, while maintaining computational efficiency.
Our work makes several key contributions to the field of multi-agent navigation. We develop a distributed navigation policy capable of operating in stochastic environments, incorporating an LSTM-based architecture to address information loss in POMDPs. The proposed physics-integrated training platform enables safe and efficient RL validation, complemented by a novel reward function that effectively resolves suboptimal action selection in complex scenarios. Comprehensive evaluations in diverse environments demonstrate the framework’s effectiveness, robustness, and real-world applicability.
In the future, we will identify multiple promising directions for future research. These include extending the framework to handle dynamic traffic scenarios and unmanned aerial vehicle (UAV) applications, refining the reward function to address complex edge cases such as U-shaped obstacles, and enhancing policy robustness in adversarial environments. The proposed framework’s adaptability and performance suggest strong potential for broader applications in autonomous navigation systems.